# ExGRG: Explicitly-Generated Relation Graph for Self-Supervised Representation Learning

### v1

@@ B.7. Selected Hyperparameters
Table 6. Selected hyperparameters utilized for each dataset.
Hyperparameter tuning is conducted to identify optimal parameters for each utilized dataset. The chosen hyperparameters are outlined in Table 6, corresponding to the models whose performance is detailed in Table 1. Due to constraints on the GPU with a 12 GB memory, a limited number of nodes is randomly selected at each iteration to construct the relation graph ${\bm{G}}$; this quantity is denoted as the mini-batch size $N$ (${\bm{G}}$ B.S.). The optimizer algorithm utilized for end-to-end model training, including AdamW (Loshchilov and Hutter, 2017) or Adam (Kingma and Ba, 2014) algorithms, is denoted as Optim. The learning rate is represented by L.R., and #Iters. signifies the number of iterations used for training the SSL encoder and expander without utilizing labels. No learning scheduler is employed. In the downstream evaluation, mirroring methods like BGRL (Thakoor et al., 2021), the expander is replaced with an $L_{2}$-regularized logistic regression classifier, while the encoder weights remain fixed.
The coefficients for the variance, covariance, and invariance loss terms introduced in Eq. 1 and 2 are denoted by $\alpha$, $\beta$, and $\gamma$, respectively. We employed various activation functions in the SSL encoder and expander, including PReLU (He et al., 2015), ReLU, and ELU (Clevert et al., 2015). Additionally, for normalization techniques, we utilized BN (Batch Normalization) (Ioffe and Szegedy, 2015) or None (no normalization). The parameters $[f_{\theta}]$ Act., Norm., #Lay., #Hid., and $D_{H}$ denote the SSL encoder’s activation function, normalization technique, number of GCN layers, number of features in hidden layers, and representation dimension, i.e., the last GCN output dimension, respectively. The same set of hyperparameters is reported for the expander $g_{\phi}$ modeled with an MLP.
@@ B.8. Hyperparameter Analysis
Figure 10. Impact of simultaneously altering $D_{H}$ and $D_{Z}$ in the particular case where $D_{H}=D_{Z}$ on our metrics and downstream accuracy for Amazon Computers.
@@
Figure 11. Impact of altering $N$, ${\bm{G}}$ mini-batch size, on our metrics and downstream accuracy for Amazon Computers.
@@
Figure 12. Impact of altering $[{\bm{G}}^{k}]~{}k$ on our metrics and downstream accuracy for Amazon Computers.
@@
Figure 13. Impact of altering $[{\bm{G}}^{O}]~{}K$ indicating the number of prototypes on our metrics and downstream accuracy for Amazon Computers.
@@
Figure 14. Impact of altering $\alpha_{1}[\mathcal{L}_{O}]$ indicating the coefficient for optimal transport alignment term on our metrics and downstream accuracy for Amazon Computers.

### v2

@@ B.7 Selected Hyperparameters
Hyperparameter tuning is conducted to identify optimal parameters for each utilized dataset. The chosen hyperparameters are outlined in Table 6, corresponding to the models whose performance is detailed in Table 1. Due to constraints on the GPU with a 12 GB memory, a limited number of nodes is randomly selected at each iteration to construct the relation graph ${\bm{G}}$; this quantity is denoted as the mini-batch size $N$ (${\bm{G}}$ B.S.). The optimizer algorithm utilized for end-to-end model training, including AdamW (Loshchilov & Hutter, 2017) or Adam (Kingma & Ba, 2014) algorithms, is denoted as Optim. The learning rate is represented by L.R., and #Iters. signifies the number of iterations used for training the SSL encoder and expander without utilizing labels. No learning scheduler is employed. In the downstream evaluation, mirroring methods like BGRL (Thakoor et al., 2021), the expander is replaced with an $L_{2}$-regularized logistic regression classifier, while the encoder weights remain fixed.
The coefficients for the variance, covariance, and invariance loss terms introduced in Eq. 1 and 2 are denoted by $\alpha$, $\beta$, and $\gamma$, respectively. We employed various activation functions in the SSL encoder and expander, including PReLU (He et al., 2015), ReLU, and ELU (Clevert et al., 2015). Additionally, for normalization techniques, we utilized BN (Batch Normalization) (Ioffe & Szegedy, 2015) or None (no normalization). The parameters $[f_{\theta}]$ Act., Norm., #Lay., #Hid., and $D_{H}$ denote the SSL encoder’s activation function, normalization technique, number of GCN layers, number of features in hidden layers, and representation dimension, i.e., the last GCN output dimension, respectively. The same set of hyperparameters is reported for the expander $g_{\phi}$ modeled with an MLP.
@@ B.8 Hyperparameter Analysis
Figure 10: Impact of simultaneously altering $D_{H}$ and $D_{Z}$ in the particular case where $D_{H}=D_{Z}$ on our metrics and downstream accuracy for Amazon Computers.
@@
Figure 11: Impact of altering $N$, ${\bm{G}}$ mini-batch size, on our metrics and downstream accuracy for Amazon Computers.
@@
Figure 12: Impact of altering $[{\bm{G}}^{k}]~{}k$ on our metrics and downstream accuracy for Amazon Computers.
@@
Figure 13: Impact of altering $[{\bm{G}}^{O}]~{}K$ indicating the number of prototypes on our metrics and downstream accuracy for Amazon Computers.
@@
Figure 14: Impact of altering $\alpha_{1}[\mathcal{L}_{O}]$ indicating the coefficient for optimal transport alignment term on our metrics and downstream accuracy for Amazon Computers.